wes fund ฯhet around 32% studerns ,hat were involved with study Compute
Answer #1

4) The stem and leaf plot is read as 5 | 2 = 52%

The minimum value = 1%

The maximum value = 92%

Median is the central value of the ordered data.

Here there are even a number of data points. So the median is the average of two central values.

Median = (32 + 34)/2 = 33%

The first quartile is the median of the lower half of the data.

Q1 = 18%

The third quartile is the median of the upper half of the data.

Q3 = 52%

Mean is the average of values.

\small \overline{x} = \frac{\sum x}{n} where n is the number of the data points.

\small \Rightarrow \overline{x} = \frac{1802}{50} = 36.04

Mean = 36.04%

We can observe that there are more data points with low percentages than high percentages which imply the data is skewed to the right.

Also, we can observe that median(33%) is lesser that mean(36.04%) i.e \small M < \overline{x}

The interquartile range (IQR) is defined as the difference between Q3 and Q1

IQR = Q3 - Q1 = 52 - 18 = 34%

Boundaries for determining outliers are Q1 - IQR, Q3 + IQR

\small \Rightarrow Boundaries are -16, 86

Yes there is an outlier and it is 92%